# LEO-Mini: An Efficient Multimodal Large Language Model using Conditional Token Reduction and Mixture of Multi-Modal Experts

### v1

@@ Abstract
Figure 1: Performance overview of LEO-Mini-Llama3-8B. 1(a): Comparison between LEO-Mini (64 tokens) and existing token reduction MLLMs [52, 31, 68, 63, 20], where LEO-Mini outperforms all models. 1(b): Comparison with other MLLMs [54, 57, 40, 34] on 11 vision-language tasks [14, 41, 29, 22, 46, 64, 26, 55, 49, 48, 42] using Llama3-8B [44] as the LLM. LEO-Mini achieves competitive performance while using only 64 visual tokens
@@ 1 Introduction
The development of multi-modal large language models (MLLMs) [1, 54, 40, 68, 8, 45] has been significantly advanced by aligning vision models [12, 43, 28] with large-scale pre-trained language models (LLMs) [44, 6]. MLLMs, such as the LLaVA [40, 39], BLIP [8], and InternVL [5], embed image patches into visual tokens through a vision expert [50, 27, 12]. Then, those visual tokens are input into the LLM for reasoning. This has led to strong performance in image and video understanding tasks [14, 41, 29, 22, 46, 64, 32, 26], bridging the gap between vision and language models.
However, the substantial computational requirement of MLLMs presents a significant challenge to their efficiency. In MLLMs, the LLM predominantly drives computational costs, as the vision expert is smaller in comparison. For example, the commonly used vision expert, CLIP-L [50], has 0.3 billion parameters, whereas LLMs such as LLaMA [44] or Vicuna [6], have 7–8 billion and 13 billion parameters, respectively. While the vision expert is relatively lightweight, its output, i.e., the visual tokens, which is fed into the LLM along with text instruction tokens, significantly increases the computational overhead. For instance, CLIP-L [50] encodes a single image into $24\times 24=576$ visual tokens, whereas textual instructions typically consist of fewer than 100 tokens. And this becomes more challenging in high-resolution image understanding [5, 1] or video understanding [30, 62, 17], which requires either more visual tokens per image or processing multiple images.
Reducing the number of visual tokens can thus be an effective strategy for enhancing the efficiency of MLLMs, either through training an efficient compressor [68, 31] or by using a training-free summarizer [3, 67, 59]. In this work, we focus on training-based methods. Recent training-based approaches [68, 31, 33, 52] reduce the number of visual tokens by selecting only the most informative ones [3], rather than using all visual tokens. Notably, LLaVA-Mini [68] achieves comparable performance to the full LLaVA-1.5 [39], while using only a single visual token. However, aggressively reducing visual tokens may result in the loss of essential visual information, potentially degrading the model’s performance.
@@
Effectiveness. For a better understanding of visual features and improved reasoning ability, LEO-Mini incorporates MMoE, a novel mixture of multi-modal experts module consisting of MMoE-LLM and MMoE-Vision. Instead of conducting a full finetuning of the entire model after the pretraining, MMoE-LLM employs a mixture of LoRA experts [19, 9, 61, 56] with a novel router and a general expert. In contrast to previous work [19, 9, 61, 56] whose routers only take the hidden state as input to switch between experts, our router takes the text tokens and the visual tokens as additional input. This facilitates more effective switching between different LoRA experts. The general expert is continuously activated to learn general knowledge. To extract more informative visual features, MMoE-Vision incorporates multiple vision experts [50, 27, 12], each trained on data from different domains. This boosts the model’s ability to understand visual information, leading to improved performance on vision-language tasks while maintaining minimal computational overhead, as both the vision experts and LoRA experts are substantially smaller than the LLM.
@@
  • To enhance reasoning ability without huge computational overhead, MMoE-LLM employs a novel router taking visual and text tokens as additional input for better switching between different experts, with a general expert for learning general knowledge. MMoE-Vision incorporates multiple vision experts for rich visual feature extraction.
  • We demonstrate the effectiveness and improved efficiency of LEO-Mini on various vision-language tasks [14, 41, 29, 22, 46, 64, 32, 26, 15, 16, 55, 49, 48], as illustrated in Fig. 1.
@@ 2 Related Work
Multi-modal large language models (MLLMs). Advancements in LLMs have fueled significant progress in MLLMs, enabling effective cross-modal reasoning through modality fusion and instruction tuning [8, 37, 39, 35, 40]. Early MLLMs struggled with complex visual understanding due to input resolution limits and the inefficiencies of single vision encoders. To address this, recent research has enhanced visual experts [5, 66], incorporated higher-resolution inputs [34, 47], and explored mixtures of vision experts [1, 45, 25, 54, 69, 11]. Despite the success of these methods, a major challenge remains: the efficiency of MLLMs, as these approaches increase the number of visual tokens, leading to higher computational costs and scalability constraints.
@@
Compressing visual tokens for MLLMs. The efficiency of MLLMs is constrained by the LLM backbone’s context length, as high-resolution images generate numerous vision tokens that quickly consume available space and increase computational cost. To address this challenge, recent methods have focused on reducing the number of visual tokens through both training-free [3, 21, 58] and training-based [68, 3, 33, 52] token reduction strategies. Focusing on training-based approaches, some models aggregate tokens based on visual feature similarities [52] or high-low resolution similarities [31], while others use attention distillation [63]. MQT-LLaVA [20] employs a query transformer to process a random subset of latent query tokens per step. However, direct compression may lead to information loss. LLaMA-VID [33] integrates text tokens as contextual information and applies average pooling for efficient token reduction. More recently, LLaVA-Mini [68] mitigates this by combining query-based reduction with prefusion of visual and text tokens.
Figure 2: The overview of the proposed LEO-Mini. LEO-Mini is designed with MMoE (Sec. 3.3) to enhance visual comprehension and CoTR (Sec. 3.2) to reduce the number of visual tokens for efficiency.
@@
Existing token reduction approaches [68, 33, 20] tend to select tokens by a learnable query or saliency maps. To the best of our knowledge, we propose the first token reduction method that uses text tokens and visual tokens from other visual experts as context to perform attention and reduction over the current vision expert’s tokens.
@@
Mixture of Experts (MoE) in LLMs. MoE is a model design that exploits multiple sparse experts to process different parts of the input space [23]. Early works [10, 13] demonstrated that sparse expert activation improves scalability and computational efficiency. Existing MoE-based LLMs [24, 7] typically incorporate MoE by replacing standard feed-forward networks with MoE layers, where each token is routed to a small subset of experts. More recent research [38, 36, 65, 60] explores integrating MoE with LoRA to further reduce the parameter overhead of traditional MoE models. These methods make use of LoRA’s ability to fine-tune only a small subset of parameters [61, 9, 4], enabling efficient expert selection and dynamic task adaptation.
@@ 3 Methodology
In Sec. 3.1, we introduce the overall framework of LEO-Mini. In Sec. 3.2, we present our proposed token reduction module, CoTR, which consolidates a large number of visual tokens into a smaller, more informative set. Finally, in Sec. 3.3, we introduce our mixture of multi-modal experts module, MMoE, which is designed to enhance the efficiency of fine-tuning MLLMs while preserving their strong performance.
@@ 3.1 Architecture
The overall architecture of LEO-Mini is presented in Fig. 2. It follows the general design (vision expert-projector-LLM) of existing MLLMs [54, 1, 68], while incorporating a mixture of multi-modal experts and a visual feature compression module.
@@
Visual feature compression is then applied to the group of visual tokens $\left\{I_{i}\right\}_{i\in[m]}$. First, as the visual tokens generated by different vision experts have different lengths, i.e., $\{N_{i}^{V}\}_{i\in[m]}$, the CoTR module (Sec. 3.2) projects them into a group of consolidated visual tokens $\left\{\bar{I}_{i}\in\mathcal{R}^{N^{V}\times d^{V}_{i}}\right\}_{i\in[m]}$ with the same length of $N^{V}$. $N^{V}$ is much smaller than $\sum_{i\in[m]}N_{i}^{V}$, significantly reducing the number of visual tokens for efficiency. Then, the consolidated visual tokens are concatenated channel-wise to form the concatenated visual tokens $\bar{I}\in\mathcal{R}^{N^{V}\times d^{V}}$, where $d^{V}=\sum_{i\in[m]}d_{i}^{V}$.
@@ 3.2 Conditional Token Reduction (CoTR)
The CoTR module, illustrated in Fig. 3, is a query-based module that takes a group of visual tokens $\{I_{i}\}_{i\in[m]}$, text tokens $T$, and a group of query tokens $\{Q_{i}\in\mathcal{R}^{N^{V}\times d_{i}^{V}}\}_{i\in[m]}$ as input, and outputs a group of consolidated visual tokens $\{\bar{I}_{i}\}_{i\in[m]}$. Specifically, for the visual tokens $I_{i}$ generated by the $i$-th vision expert, the CoTR module computes an attention score $\alpha_{i}\in\mathcal{R}^{N^{V}\times N^{V}_{i}}$ using the query token $Q_{i}\in\mathcal{R}^{N^{V}\times d_{i}^{V}}$, text tokens $T$, the visual tokens from other vision experts $\left\{I_{j}\right\}_{j\in[m]\setminus\{i\}}$, and the visual token $I_{i}$, as
@@ 3.3 Mixture of Multi-modal Experts (MMoE)
Effective visual comprehension (MMoE-Vision). As described in Sec. 3.1, drawing inspiration from previous work [54, 1], visual tokens are generated by multiple vision experts, each extracting informative features from different perspectives to enrich visual understanding.
Effective reasoning ability (MMoE-LLM). To ensure efficient training with minimal computational overhead, we introduce MMoE for LLM tuning, following the mixture of LoRA experts [19], as shown in Fig. 4. The vanilla mixture of LoRA experts consists of a set of experts [9, 56], i.e., $\{f^{E}_{i}(\cdot)\}_{i\in[E]}$, and a routing network $f^{\textsc{routing}}(\cdot)$ that outputs the routing probability $R\in\mathcal{R}^{E}$ taking the hidden state of the current layer as input, where $E$ is the number of experts. Then, based on the routing probability, only the top-$k$ experts will be activated.

### v2

@@ Abstract
LEO-Mini: An Efficient Multimodal Large Language Model using Conditional Token Reduction and Mixture of Multi-Modal Experts
Yimu Wangthanks: Equal contribution, Mozhgan Nasr Azadani, Sean Sedwards, Krzysztof Czarnecki University of Waterloo, Canada {yimu.wang,mnasraza,sean.sedwards,k2czarne}@uwaterloo.ca
@@
Figure 1: Performance overview of LEO-Mini-Llama3-8B. 1(a): Comparison between LEO-Mini (64 tokens) and existing token reduction MLLMs , where LEO-Mini outperforms all models. 1(b): Comparison with other MLLMs on $1$ vision-language tasks using Llama3-8B as the LLM. LEO-Mini achieves competitive performance while using only 64 visual tokens.
@@ 1 Introduction
The development of multi-modal large language models (MLLMs) Azadani et al. (2025); Shi et al. (2024); Liu et al. (2024c); Zhang et al. (2025); Dai et al. (2023); Lu et al. (2024); Wang et al. (2025a) has been significantly advanced by aligning vision models Fang et al. (2024c); Liu et al. (2022); Lee et al. (2023) with large-scale pre-trained language models (LLMs) Llama Team (2024); Chiang et al. (2023). MLLMs, such as the LLaVA Liu et al. (2024c, b), BLIP Dai et al. (2023), and InternVL Chen et al. (2024c), embed image patches into visual tokens through a vision expert Radford et al. (2021); Kirillov et al. (2023); Fang et al. (2024c). Then, those visual tokens are input into the LLM for reasoning. This has led to strong performance in image and video understanding tasks Fu et al. (2024); Liu et al. (2024d); Li et al. (2024a); Hudson and Manning (2019); Lu et al. (2022); Yue et al. (2024); Li et al. (2023); Kembhavi et al. (2016), bridging the gap between vision and language models.
However, the substantial computational requirement of MLLMs presents a significant challenge to their efficiency. In MLLMs, the LLM predominantly drives computational costs, as the vision expert is smaller in comparison. For example, the commonly used vision expert, CLIP-L Radford et al. (2021), has 0.3 billion parameters, whereas LLMs such as LLaMA Llama Team (2024) or Vicuna Chiang et al. (2023), have 7–8 billion and 13 billion parameters, respectively. While the vision expert is relatively lightweight, its output, i.e., the visual tokens, which is fed into the LLM along with text instruction tokens, significantly increases the computational overhead. For instance, CLIP-L Radford et al. (2021) encodes a single image into $24\times 24=576$ visual tokens, whereas textual instructions typically consist of fewer than 100 tokens. And this becomes more challenging in high-resolution image understanding Chen et al. (2024c); Azadani et al. (2025) or video understanding Li et al. (2024b); Xu et al. (2016); Heilbron et al. (2015), which requires either more visual tokens per image or processing multiple images.
Reducing the number of visual tokens can thus be an effective strategy for enhancing the efficiency of MLLMs, either through training an efficient compressor Zhang et al. (2025); Li et al. (2024c) or by using a training-free summarizer Chen et al. (2024a); Zhang et al. (2024); Wen et al. (2025a). In this work, we focus on training-based methods. Recent training-based approaches Zhang et al. (2025); Li et al. (2024c, d); Shang et al. (2024) reduce the number of visual tokens by selecting only the most informative ones Chen et al. (2024a), rather than using all visual tokens. Notably, LLaVA-Mini Zhang et al. (2025) achieves comparable performance to the full LLaVA-1.5 Liu et al. (2024b), while using only a single visual token. However, aggressively reducing visual tokens may result in the loss of essential visual information, potentially degrading the model’s performance.
@@
Effectiveness. For a better understanding of visual features and improved reasoning ability, LEO-Mini incorporates MMoE, a novel mixture of multi-modal experts module consisting of MMoE-LLM and MMoE-Vision. Instead of conducting a full finetuning of the entire model after the pretraining, MMoE-LLM employs a mixture of LoRA experts Hu et al. (2022); Dou et al. (2024); Wu et al. (2023); Tian et al. (2024) with a novel router and a general expert. In contrast to previous work Hu et al. (2022); Dou et al. (2024); Wu et al. (2023); Tian et al. (2024) whose routers only take the hidden state as input to switch between experts, our router takes the text tokens and the visual tokens as additional input. This facilitates more effective switching between different LoRA experts. The general expert is continuously activated to learn general knowledge. To extract more informative visual features, MMoE-Vision incorporates multiple vision experts Radford et al. (2021); Kirillov et al. (2023); Fang et al. (2024c), each trained on data from different domains. This boosts the model’s ability to understand visual information, leading to improved performance on vision-language tasks while maintaining minimal computational overhead, as both the vision experts and LoRA experts are substantially smaller than the LLM.
@@
  • MMoE-LLM employs a novel router taking visual and text tokens as additional input for better switching between different experts, with a general expert for learning general knowledge. MMoE-Vision incorporates multiple vision experts for rich visual feature extraction.
  • We demonstrate the effectiveness and improved efficiency of LEO-Mini on various vision-language tasks, as illustrated in Figure˜1.
@@ 2 Related Work
Multi-modal large language models (MLLMs). As a fundamental problem in multimodal learning Fang et al. (2022, 2023, 2024b, 2024a, 2025); Wang et al. (2021, 2020a, 2023, 2020b, 2025b); Liu et al. (2025); Wang and Shi (2023); Li et al. (2023); Liu et al. (2024d), visual understanding has been revolutionized by recent developments in large language models (LLMs). Advancements in LLMs have fueled significant progress in MLLMs, enabling effective cross-modal reasoning through modality fusion and instruction following Dai et al. (2023); Lin et al. (2024b); Liu et al. (2024b); Li et al. (2024f); Liu et al. (2024c). Early MLLMs struggled with complex visual understanding due to input resolution limits and the inefficiencies of single vision encoders. To address this, recent research has enhanced visual experts Chen et al. (2024c); Zhai et al. (2023), incorporated higher-resolution inputs Li et al. (2024e); Luo et al. (2024), and explored mixtures of vision experts Azadani et al. (2025); Lu et al. (2024); Kar et al. (2024); Shi et al. (2024); Zong et al. (2024); Fan et al. (2024). Despite the success of these methods, a major challenge remains: the efficiency of MLLMs, as these approaches increase the number of visual tokens, leading to higher computational costs and scalability constraints.
@@
Compressing visual tokens for MLLMs. The efficiency of MLLMs is constrained by the LLM backbone’s context length, as high-resolution images generate numerous vision tokens that quickly consume available space and increase computational cost. To address this challenge, recent methods have focused on reducing the number of visual tokens through both training-free Chen et al. (2024a); Huang et al. (2024); Wen et al. (2025b) and training-based Zhang et al. (2025); Chen et al. (2024a); Li et al. (2024d); Shang et al. (2024) token reduction strategies. Focusing on training-based approaches, some models aggregate tokens based on visual feature similarities Shang et al. (2024) or high-low resolution similarities Li et al. (2024c), while others use attention distillation Ye et al. (2024). MQT-LLaVA Hu et al. (2024) employs a query transformer to process a random subset of latent query tokens per step. However, direct compression may lead to information loss. LLaMA-VID Li et al. (2024d) integrates text tokens as contextual information and applies average pooling for efficient token reduction. More recently, LLaVA-Mini Zhang et al. (2025) mitigates this by combining query-based reduction with prefusion of visual and text tokens.
@@
Existing token reduction approaches Zhang et al. (2025); Li et al. (2024d); Hu et al. (2024) tend to select tokens by a learnable query or saliency maps. To the best of our knowledge, we propose the first token reduction method that uses text tokens and visual tokens from other visual experts as context to perform attention and reduction over the current vision expert’s tokens.
@@
Mixture of Experts (MoE) in LLMs. MoE is a model design that exploits multiple sparse experts to process different parts of the input space Jacobs et al. (1991). Early works Du et al. (2022); Fedus et al. (2022) demonstrated that sparse expert activation improves scalability and computational efficiency. Existing MoE-based LLMs Jiang et al. (2024); Dai et al. (2024) typically incorporate MoE by replacing standard feed-forward networks with MoE layers, where each token is routed to a small subset of experts. More recent research Liu et al. (2024a); Lin et al. (2024a); Zadouri et al. (2023); Wu et al. (2024) explores integrating MoE with LoRA to further reduce the parameter overhead of traditional MoE models. These methods make use of LoRA’s ability to fine-tune only a small subset of parameters Wu et al. (2023); Dou et al. (2024); Chen et al. (2024b), enabling efficient expert selection and dynamic task adaptation.
@@ 3 Methodology
In Section˜3.1, we introduce the overall framework of LEO-Mini. In Section˜3.2, we present our proposed token reduction module, CoTR, which consolidates a large number of visual tokens into a smaller, more informative set. Finally, in Section˜3.3, we introduce our mixture of multi-modal experts module, MMoE, which is designed to enhance the efficiency of fine-tuning MLLMs while preserving their strong performance.
Figure 2: The overview of the proposed LEO-Mini. LEO-Mini is designed with MMoE (Section˜3.3) to enhance visual comprehension and CoTR (Section˜3.2) to reduce the number of visual tokens for efficiency.
@@ 3.1 Architecture
The overall architecture of LEO-Mini is presented in Figure˜2. It follows the general design (vision expert-projector-LLM) of existing MLLMs Shi et al. (2024); Azadani et al. (2025); Zhang et al. (2025), while incorporating a mixture of multi-modal experts and a visual feature compression module.
@@
Visual feature compression is then applied to the group of visual tokens $\left\{I_{i}\right\}_{i\in[m]}$. First, as the visual tokens generated by different vision experts have different lengths, i.e., $\{N_{i}^{V}\}_{i\in[m]}$, the CoTR module (Section˜3.2) projects them into a group of consolidated visual tokens $\left\{\bar{I}_{i}\in\mathcal{R}^{N^{V}\times d^{V}_{i}}\right\}_{i\in[m]}$ with the same length of $N^{V}$. $N^{V}$ is much smaller than $\sum_{i\in[m]}N_{i}^{V}$, significantly reducing the number of visual tokens for efficiency. Then, the consolidated visual tokens are concatenated channel-wise to form the concatenated visual tokens $\bar{I}\in\mathcal{R}^{N^{V}\times d^{V}}$, where $d^{V}=\sum_{i\in[m]}d_{i}^{V}$.
@@ 3.2 Conditional Token Reduction (CoTR)
The CoTR module, illustrated in Figure˜3, is a query-based module that takes a group of visual tokens $\{I_{i}\}_{i\in[m]}$, text tokens $T$, and a group of query tokens $\{Q_{i}\in\mathcal{R}^{N^{V}\times d_{i}^{V}}\}_{i\in[m]}$ as input, and outputs a group of consolidated visual tokens $\{\bar{I}_{i}\}_{i\in[m]}$. Specifically, for the visual tokens $I_{i}$ generated by the $i$-th vision expert, the CoTR module computes an attention score $\alpha_{i}\in\mathcal{R}^{N^{V}\times N^{V}_{i}}$ using the query token $Q_{i}\in\mathcal{R}^{N^{V}\times d_{i}^{V}}$, text tokens $T$, the visual tokens from other vision experts $\left\{I_{j}\right\}_{j\in[m]\setminus\{i\}}$, and the visual token $I_{i}$, as
@@
Model # of General OCR Knowledge
[…]
VoCo-LLaMA Ye et al. (2024) 1 1323.3 58.8 53.7 57.0 - - 65.4 81.4 -
LLaMA-VID Li et al. (2024d) 2 - - - 55.5 - 49.0 67.7 83.1 -
PruMerge Shang et al. (2024) 32 1350.3 60.9 - - - 56.0 72.0 76.3 -
MQT-LLaVA Hu et al. (2024) 64 1464.3 63.5 - 60.0 51.5 - 67.0 83.6 34.4
Token-Packer Li et al. (2024c) 64 - 64.1 - 61.1 50.7 - - 86.3 -
LLaVA-Mini Zhang et al. (2025) 64 1476.8 67.5 60.2 61.8 58.5 59.1 69.7 85.3 -
LEO-Mini-Vicuna-7B 64 1542.6 67.8 73.2 64.0 51.4 70.1 73.3 90.0 34.1
LEO-Mini-Llama-8B 64 1583.0 77.0 75.8 64.5 69.3 75.1 84.5 90.3 38.8
Model MMEP SEEDI GQA SQA MMMU POPE AI2D TextVQA ChartQA OCRBench
LEO-Mini (64 tokens) 1583 75.8 64.5 84.5 38.8 90.3 75.7 75.1 80.5 62.4
w/ 1 visual token 1565 74.6 64.2 83.6 37.7 89.0 74.5 73.5 80.0 61.7
w/ 16 visual token 1577 75.4 64.4 84.3 38.2 90.1 75.4 74.2 80.2 62.6
w/ 256 visual token 1584 76.1 64.2 85.5 39.0 90.7 75.5 75.5 80.5 63.0
w/ 1.8m SFT data 1548 76.6 64.1 85.9 39.4 90.3 77.9 75.7 80.9 63.1
@@ 3.3 Mixture of Multi-modal Experts (MMoE)
Effective visual comprehension (MMoE-Vision). As described in Section˜3.1, drawing inspiration from previous work Shi et al. (2024); Azadani et al. (2025), visual tokens are generated by multiple vision experts, each extracting informative features from different perspectives to enrich visual understanding.
Effective reasoning ability (MMoE-LLM). To ensure efficient training with minimal computational overhead, we introduce MMoE for LLM tuning, following the mixture of LoRA experts Hu et al. (2022), as shown in Figure˜4. The vanilla mixture of LoRA experts consists of a set of experts Dou et al. (2024); Tian et al. (2024), i.e., $\{f^{E}_{i}(\cdot)\}_{i\in[E]}$, and a routing network $f^{\textsc{routing}}(\cdot)$ that outputs the routing probability $R\in\mathcal{R}^{E}$ taking the hidden state as input, where $E$ is the number of experts. Then, based on the routing probability, only the top-$1$ expert will be activated.